Pham can work as many or as few hours as she wants at the college bookstore for $9 per hour. But due to her hectic schedule, she has just 18 hours per week that she can spend working at either the bookstore or at other potential jobs. One potential job, at a café, will pay her $12 per hour for up to 7 hours per week. She has another job offer at a garage that will pay her $10 an hour for up to 5 hours per week. And she has a potential job at a daycare center that will pay her $8.50 per hour for as many hours as she can work. Instructions: Enter your answer as a whole number. If her goal is to maximize the amount of money she can make each week, how many hours will she work at the bookstore? hour(s)
Answer #1

Pham can earn TR1 = $9*18= $162 at the college bookstore.
In cafe she can earn TR2 = $12*7= $84
In garage she can earn TR3 = $10*5 = $50.
In daycare center she can earn TR4 = $8.50*18=$153


Fir revenue maximization she should work 18 hours per week at the bookstore to earn $162.